### v2

# arXiv:2511.10788v2[cs.AI] 01 Mar 2026

Abstract
Recent advances in large language models (LLMs) have made reasoning a central benchmark for evaluating intelligence. While prior surveys focus on efficiency by examining how to shorten reasoning chains or reduce computation, this view overlooks a fundamental challenge: current LLMs apply uniform reasoning strategies regardless of task complexity, generating long traces for trivial problems while failing to extend reasoning for difficult tasks. This survey reframes reasoning through the lens of adaptivity: the capability to allocate reasoning effort based on input characteristics such as difficulty and uncertainty. We make three contributions. First, we formalize deductive, inductive, and abductive reasoning within the LLM context, connecting these classical cognitive paradigms with their algorithmic realizations. Second, we formalize adaptive reasoning as a control-augmented policy optimization problem balancing task performance with computational cost, distinguishing learned policies from inference-time control mechanisms. Third, we propose a systematic taxonomy organizing existing methods into training-based approaches that internalize adaptivity through reinforcement learning, supervised fine-tuning, and learned controllers, and training-free approaches that achieve adaptivity through prompt conditioning, feedback-driven halting, and modular composition. This framework clarifies how different mechanisms realize adaptive reasoning in practice and enables systematic comparison across diverse strategies. We conclude by identifying open challenges in self-evaluation, meta-reasoning, and human-aligned reasoning control.
From Efficiency to Adaptivity: A Deeper Look at Adaptive Reasoning in Large Language Models
Chao Wu University at Buffalo cwu64@buffalo.edu          Baoheng Li University at Buffalo baohengl@buffalo.edu
Mingchen Gao University at Buffalo mgao8@buffalo.edu          Yu Tian University of Central Florida yu.tian2@ucf.edu          Zhenyi Wang University of Central Florida zhenyi.wang@ucf.efu
1 Introduction
Reasoning is a core aspect of intelligence, allowing humans to draw inferences from evidence, rules, and prior knowledge to make systematic decisions. In the context of large language models (LLMs), reasoning has become a central benchmark of progress, not only for solving complex tasks such as mathematics, programming, and scientific discovery, but also for demonstrating abilities beyond pattern recognition. As LLMs scale in capability, understanding how they reason and how to control their reasoning processes has emerged as a fundamental research question with both theoretical and practical implications.
Existing surveys on LLM reasoning have largely focused on efficiency: how to shorten reasoning chains, avoid overthinking, and reduce computational cost. While efficiency is an important practical concern, it does not fully capture the broader challenge. Current LLMs often adopt a one-size-fits-all reasoning strategy, generating long reasoning traces even for trivial problems while failing to extend reasoning when faced with difficult tasks. What is missing is not merely efficiency, but adaptivity, which is the ability to flexibly adjust reasoning strategies based on context, difficulty, and uncertainty. We define adaptive reasoning in LLMs as the capability to allocate reasoning effort based on input characteristics such as difficulty and uncertainty. This adaptivity can be realized through learned policies that internalize allocation strategies during training, or through dynamic control mechanisms applied at inference time. Unlike efficiency-centric approaches that uniformly constrain reasoning length, adaptive reasoning emphasizes input-dependent resource allocation: answering easy questions quickly while allocating deliberate reasoning effort to harder or unfamiliar problems, thereby mirroring the natural scaling of human cognitive effort with task demands.
To realize adaptive reasoning, LLMs must flexibly employ different reasoning strategies depending on task demands. Reasoning in LLMs encompasses diverse cognitive patterns, among which three classical paradigms—inductive, deductive, and abductive reasoning—provide a useful conceptual lens. Inductive reasoning generalizes from observed examples, as seen in in-context learning Brown et al. (2020); deductive reasoning applies explicit logical rules to derive valid conclusions, underlying structured chain-of-thought Wei et al. (2022) and program-of-thought Chen et al. (2022) frameworks; abductive reasoning seeks plausible explanations for observed outcomes. However, rather than organizing methods by reasoning type, this survey focuses on how adaptivity is implemented in practice. We propose a systematic taxonomy distinguishing training-based approaches, which internalize adaptive allocation strategies through reinforcement learning, supervised fine-tuning, and learned controllers, from training-free approaches, which achieve adaptivity through prompt conditioning, feedback-driven halting, and modular composition at inference time. Our contributions are threefold:
  • We reframe LLM reasoning research through the lens of adaptive reasoning, moving beyond efficiency-centric approaches to emphasize input-dependent resource allocation. We define adaptivity as the capability to allocate reasoning effort based on task difficulty and uncertainty, realized through either learned policies or dynamic inference control.
  • We formalize deductive, inductive, and abductive reasoning within the LLM context, providing clear definitions that connect these classical cognitive paradigms with their algorithmic realizations in contemporary LLM systems.
  • We propose a systematic taxonomy organizing existing methods by implementation mechanism, distinguishing training-based approaches that learn adaptive policies from training-free approaches that apply dynamic control at inference and enabling systematic comparison across diverse adaptive reasoning strategies.
2 Background and Foundations
2.1 What is reasoning in LLMs?
Reasoning is a cornerstone of human intelligence, enabling inference and decision-making from evidence and rules. Following Huang and Chang (2023) and Sun et al. (2025), reasoning can be viewed as a logical and systematic process that uses prior knowledge and contextual evidence to arrive at valid conclusions or hypotheses. Within the context of LLMs, reasoning refers to the model’s capacity to generate and verify intermediate representations that connect an input prompt $\mathbf{x}$ to an output $y$ through a structured inference procedure.
We formalize reasoning from two complementary perspectives: (1) as a latent-variable conditional generation process (§2.1.1), and (2) as a resource-bounded optimization problem (§2.1.2). These formalizations provide the mathematical foundation for understanding adaptive reasoning mechanisms.
2.1.1 Reasoning as Conditional Generation
Following Bandyopadhyay et al. (2025), A language model defines a conditional probability distribution over output sequences:
$p_{\bm{\theta}}(\mathbf{y}\mid\mathbf{x})=\prod_{n=1}^{N}p_{\bm{\theta}}(y_{n}\mid\mathbf{x},y_{<n}),$ (1)
where $\bm{\theta}$ denotes model parameters.
Extending this formulation, Phan et al. (2023) formalize reasoning as a latent-variable conditional generation process. Specifically, the model can be viewed as first generating a latent reasoning trace $\mathbf{r}$, which is an unobserved sequence of intermediate thought step, and then producing the final output $\mathbf{y}$ conditioned on both $\mathbf{x}$ and $\mathbf{r}$. This leads to the marginal likelihood:
$\log p_{\bm{\theta}}(\mathbf{y}\mid\mathbf{x})=\log\sum_{\mathbf{r}}p_{\bm{\theta}}(\mathbf{r}\mid\mathbf{x})\,p_{\bm{\theta}}(\mathbf{y}\mid\mathbf{r},\mathbf{x}).$ (2)
Under this latent-variable perspective, reasoning is formalized as marginalizing over possible reasoning trajectories that mediate between the input $\mathbf{x}$ and the final output $\mathbf{y}$. This view conceptually unifies reasoning and conditional text generation, suggesting that the “thought” process of an LLM can be understood as sampling from an implicit distribution over intermediate reasoning states.
2.1.2 Reasoning as Resource-Bounded Optimization
Following Alomrani et al. (2025), test-time reasoning can be viewed as a trade-off between computation and performance. Two equivalent optimization views are commonly considered: a constrained form ($L_{1}$) that fixes the compute budget, and a penalized form ($L_{2}$) that adaptively balances accuracy and efficiency. We adopt the $L_{2}$ formulation, which naturally aligns with the idea of adaptive reasoning:
$\max_{\mathbf{r}\in\mathcal{R}}\;\mathcal{P}(\mathbf{r},\mathbf{x})+\alpha\,\mathcal{E}(\mathbf{r},\mathbf{x}),$ (3)
where $\mathcal{P}$ measures task performance, $\mathcal{E}$ quantifies computational efficiency (e.g., the inverse of compute cost), and $\alpha$ controls their trade-off. Although $L_{2}$ removes the explicit resource constraint in $L_{1}$, it serves as its Lagrangian relaxation, implicitly preserving the same budgeted principle through the weighted efficiency term. This formulation highlights reasoning as an adaptive optimization process under limited computational resources.
2.1.3 Limitations of Efficient Reasoning
LLMs often exhibit an overthinking phenomenon, producing unnecessarily long chains of thought (CoT) for simple problems, thereby increasing compute without improving accuracy (Sui et al., 2025a; Alomrani et al., 2025). Efficiency-oriented methods attempt to limit reasoning tokens, steps, or decoding paths to reduce test-time compute. However, such static control allocates identical budgets to all inputs, regardless of difficulty or confidence, leading to redundant reasoning on easy cases and insufficient reasoning on complex ones.
Empirical studies confirm that fixed-length CoT models often waste computation on easy inputs while failing on hard ones (Sui et al., 2025a). The key limitation is that computation is treated as an external constraint rather than a self-regulated process: intelligent reasoning should dynamically decide how much to think based on uncertainty and task complexity.
To address these limitations, this survey proposes adaptive reasoning as a unified framework that (1) characterizes reasoning through broader cognitive paradigms (§2.2), and (2) formalizes adaptivity as a policy optimization problem (§2.3), distinguishing methods that internalize adaptivity through training from those that apply external control at inference.
2.2 Forms of Reasoning
We formalize three classical reasoning paradigms that LLMs may employ. Let $\mathbf{x}$ denote the input, $\mathbf{y}$ the output, and $\mathbf{r}=(r_{1},\ldots,r_{K})$ the intermediate reasoning steps. Each paradigm defines how the model constructs and evaluates reasoning trajectories.
Setup.
A reasoning process generates an intermediate trajectory $\mathbf{r}=(r_{1},\ldots,r_{K})$ that mediates between input $\mathbf{x}$ and output $\mathbf{y}$, where each step $r_{k}$ is autoregressively conditioned on $(\mathbf{x},r_{<k})$. We call a reasoning episode $R_{\bm{\theta}}(\mathbf{x})=(\mathbf{r},\mathbf{y})$ valid if it satisfies (1) Causal coherence — each step $r_{k}$ depends only on $\mathbf{x}$ and previous steps, ensuring a forward information flow; (2) Goal consistency — the reasoning trajectory $\mathbf{r}$ produces an answer $\mathbf{y}$ that faithfully and correctly fulfills the task objective, following the principles of faithful and step-wise reasoning (Lee and Hockenmaier, 2025; Lyu et al., 2023)
Definition 2.1 (Deductive Reasoning).
A reasoning process is deductive if each intermediate step $r_{k}$ logically follows from the input $\mathbf{x}$ and prior steps, and the final answer $y$ is necessarily entailed:
$\mathbf{x}\cup\{r_{1},\dots,r_{k-1}\}\models r_{k},\quad$ $\forall k\in\{1,\dots,K\},$ (4)
$\mathbf{x}\cup\mathbf{r}\models y.$
Here, the symbol $\models$” denotes logical entailment: a set of premises $A$ entails a conclusion $B$ ($A\models B$) if every model that satisfies all formulas in $A$ also satisfies $B$.
This ensures that the reasoning trajectory preserves logical validity at every step (Ling et al., 2023). In LLMs, deductive reasoning is approximated through verifiable chain-of-thought generation, where each step can be checked for logical entailment from previous statements (Ling et al., 2023; Seals and Shalin, 2024; Xia et al., 2025b).
Definition 2.2 (Inductive Reasoning).
Let $E=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{K}$ be observed examples, $\mathcal{H}$ a hypothesis space $h:\mathcal{X}\to\mathcal{Y}$, and $C:\mathcal{H}\to\mathbb{R}_{\geq 0}$ a complexity measure. An inductive reasoning process infers:
$h^{*}=\arg\min_{h\in\mathcal{H}:h(\mathbf{x}_{i})=y_{i}\,\forall i}C(h),$ (5)
the simplest hypothesis consistent with all observations.
This formalizes Occam’s Razor (Baker, 2022) that preferring hypotheses that explain data with minimal complexity. When $C(h)=-\log p(h)$, this reduces to Maximum A Posteriori (MAP) estimation under Bayesian induction. In LLMs, inductive reasoning manifests primarily in in-context learning (ICL) (Brown et al., 2020), where models extrapolate implicit rules from few-shot demonstrations to unseen cases (Ye et al., 2025). However, empirical studies show that LLMs do not always follow simplicity principles in practice (Sun and Saparov, 2025).
Definition 2.3 (Abductive Reasoning).
Given observations $O$ and background knowledge $\mathcal{B}$, abductive reasoning generates the most plausible explanatory hypothesis $H^{*}$ from a candidate space $\mathcal{H}$:
$H^{*}=\arg\max_{H\in\mathcal{H}}p_{\theta}(H\mid O,\mathcal{B})\quad\text{s.t.}\quad\mathcal{B}\cup H\vdash O,$ (6)
where $p_{\theta}(H\mid O,\mathcal{B})$ quantifies the model’s belief in the plausibility of each hypothesis given the observation and prior knowledge.
Here, the symbol $\vdash$” denotes syntactic entailment—that is, the observation $O$ can be formally derived from the combined premises $\mathcal{B}$ and $H$ according to a logical inference system.
This instantiates inference to the best explanation (Baker, 2022): among hypotheses that entail observation $O$ when combined with background knowledge $\mathcal{B}$, abduction selects the most probable one. In LLMs, abductive reasoning is realized through conditional generation where the model samples from $p_{\theta}(H\mid O)$ to produce plausible explanations. Benchmarks such as Abductive (Bhagavatula et al., 2020) and INABHYD (Sun and Saparov, 2025) evaluate this ability by measuring explanation coherence and plausibility.
2.3 Adaptive Reasoning
Formally, we characterize adaptive reasoning through a control-augmented policy $\pi_{\bm{\theta}}(\cdot|\mathbf{x};\bm{\phi}(\mathbf{x}))$ that maps inputs $\mathbf{x}$ to distributions over reasoning trajectories $\bm{r}$, where $\bm{\theta}$ denotes the base model parameters (e.g., LLM weights) and $\bm{\phi}$ the adaptive control function regulating reasoning behavior.
Adaptive reasoning optimizes the control policy to balance task performance $\mathcal{P}(\bm{r},\mathbf{x})$ and computational cost $\mathcal{C}(\bm{r},\mathbf{x})$:
$\max_{\ \bm{\phi}\in\Phi}\;\mathbb{E}_{\mathbf{x}\sim\mathcal{D},\,\bm{r}\sim\pi_{\bm{\theta}}(\cdot\,|\,\mathbf{x};\,\bm{\phi}(\mathbf{x}))}\big[\,\mathcal{P}(\bm{r},\mathbf{x})\;-\;\lambda\,\mathcal{C}(\bm{r},\mathbf{x})\,\big].$ (7)
where $\lambda$ controls the cost penalty and $\mathcal{D}$ is the data distribution. The optimization over $\bm{\phi}\in\Phi$ captures two complementary approaches to adaptive reasoning. In training-based methods3.1), $\bm{\phi}$ is learned during training, for instance, IBPO (Yu et al., 2025b) trains a neural network to predict reasoning budgets, while C3oT (Kang et al., 2025a) learns to generate short or long reasoning traces conditionally. In training-free methods3.2), $\bm{\phi}$ represents fixed inference-time control such as entropy-based halting (Laaouach, 2025) or adaptive sampling thresholds, applied without parameter updates. This unified formulation clarifies how adaptivity arises from either learned policies or inference-time heuristics, both of which modulate $\pi_{\bm{\theta}}$ to allocate reasoning effort based on input characteristics.
Unlike static efficiency methods that apply uniform compute budgets to all inputs, this framework enables input-dependent resource allocation by conditioning reasoning on $\bm{\phi}(\mathbf{x})$, which adapts to task difficulty and uncertainty. Recent work further shows that such adaptivity can emerge even without reinforcement learning: Karan and Du (2025) demonstrates that inference-time adaptive sampling alone can elicit strong reasoning behaviors from base LLMs, highlighting that reasoning adaptivity may arise naturally from sampling dynamics rather than explicit policy optimization.
Adaptive mechanisms in practice.
Adaptive mechanisms in practice. Existing work realizes adaptivity in two ways, corresponding to whether the policy $\pi_{\bm{\theta}}$ is optimized during training or fixed at inference, shown in our taxonomy(§1).
(1) Training-free adaptivity. Adaptivity can emerge without retraining the model, by dynamically regulating reasoning length or sampling during inference. Laaouach (2025) halt generation when the stepwise entropy $H_{i}=-\!\sum_{a}p_{i}(a)\log p_{i}(a)$ drops below a threshold $\theta$. Aggarwal et al. (2023) use a Dirichlet prior over samples to detect when majority confidence exceeds $C_{\text{thresh}}$, stopping further decoding. Han et al. (2025) estimates minimal token budgets per instance and predicts per-sample cost as $\text{Cost}=\frac{1}{N}\sum_{i=1}^{N}T(M(x_{i}))$. These methods achieve instance-level compute control through entropy, confidence, or budget signals.
(2) Training-based adaptivity. Other approaches learn a reasoning policy that internalizes computation allocation during optimization. Yu et al. (2025b) formulate inference-aware optimization with explicit budget constraints, where the policy is trained to maximize reward while satisfying correctness thresholds that implicitly controls compute allocation. Such training-based strategies generalize TTC to the learning phase, yielding policies that adapt depth and effort before deployment.
Key properties.  equation 7 directly encodes three key properties of adaptive reasoning:
[…]
(3) Cost-aware optimization—the term $\lambda\cdot\mathcal{C}(\mathbf{r})$ explicitly balances correctness and efficiency.
These properties are realized through training-based methods (§3.1) that optimize $\pi_{\theta}$ during training, or training-free methods (§3.2) that apply dynamic control at inference. This distinction forms the foundation for our taxonomy in §1.
3 Taxonomy of Adaptive Reasoning
Adaptive Reasoning Methods in LLMs
Training-based3.1)Training-free3.2) RL Methods (§3.1.1): IBPO Yu et al. (2025b), LCPO Aggarwal and Welleck (2025), Reasoning-efficient RL Arora and Zanette (2025), SABER Zhao et al. (2025), LSPO Chen et al. (2025a), LASER Liu et al. (2025b), SFPO Wang et al. (2025g), BRPO Qi et al. (2025), HBPO Lyu et al. (2025),BudgetThinker Wen et al. (2025),AdaReasoner Wang et al. (2025d) SFT & Distillation Methods (§3.1.2): C3oT Kang et al. (2025a), CoT-Valve Ma et al. (2025), TokenSkip Xia et al. (2025a), LS-Mixture SFT Yu et al. (2025a), Distilling 2-1 Yu et al. (2024), SoftCoT Xu et al. (2025b), SoftCoT++ Xu et al. (2025c), Skip-Thinking Chen et al. (2025b), MoR Xiong et al. (2025), LIFT-COT Chen and Li (2025), TALE Han et al. (2025), SCoT Wang et al. (2025b) Learned Controllers and Routers Methods (§3.1.3): RouteLLM Ong et al. (2025), SelfBudgeter Li et al. (2025b), ThinkSwitcher Liang et al. (2025),AdaMOE  Zeng et al. (2024),Route to Reason (Pan et al., 2025) Prompt-conditioned Methods (§3.2.1): Token complexity Lee et al. (2025), Constrained-CoT Nayab et al. (2025), Sketch-of-Thought Aytes et al. (2025), Chain-of-Draft Xu et al. (2025a), ConciseHint Tang et al. (2025), Plan-and-Budget Lin et al. (2025) Feedback-driven Methods (§3.2.2): Entropy halting Laaouach (2025), DEER Yang et al. (2025), ES-CoT Mao et al. (2025), EAGER Scalena et al. (2025), Self-Certainty Kang et al. (2025b), Sequence entropy Zhu et al. (2025); Sharma and Chopra (2025), DeepConf Fu et al. (2025), Answer convergence Liu and Wang (2025), Filler suppression Wang et al. (2025a), Speculative Search Wang et al. (2025f), Value-Guided Search Wang et al. (2025c), Adaptive-Consistency Aggarwal et al. (2023), RASC Wan et al. (2025), Fast BoN Sun et al. (2024), ST-BoN Wang et al. (2025e), DPTS Ding et al. (2025), Sample Smart Li et al. (2025a),MetaReasoner Sui et al. (2025b), Modular Methods (§3.2.3): AlphaOne Zhang et al. (2025), Bag-of-Tricks Liu et al. (2025a), L2S merging Wu et al. (2025), CoThink Fan et al. (2025),ACM Yao et al. (2025), Lang-specific merging Pipatanakul et al. (2025), Kimi k1.5 Kimi (2025)
Figure 1: Taxonomy of adaptive reasoning methods in LLMs.
3.1 Training-based Adaptive Reasoning
Training-based adaptive reasoning methods aim to endow models with the ability to regulate their reasoning depth and computational cost during training. Instead of applying external halting or sampling rules, these methods modify learning objectives or data organization so that the model internalizes a sense of when to reason longer, when to compress, and how to allocate computational resources based on instance difficulty or uncertainty. This section reviews five main families of approaches that operationalize such adaptivity at training time.
3.1.1 Reinforcement Learning Methods
Reinforcement-learning (RL)–based approaches integrate reasoning accuracy and computational efficiency into a unified optimization objective. IBPO (Yu et al., 2025b) formulates an inference-budget–constrained policy optimization problem where the policy learns to allocate larger token budgets to harder questions and smaller ones to easier instances. LCPO (Aggarwal and Welleck, 2025) directly adds a length-adherence(i.e., a reward that penalizes deviations from the target reasoning length) reward term to policy gradients, enabling the agent to match target reasoning lengths and revealing the emergence of short-reasoning models that preserve logical fidelity. (Arora and Zanette, 2025) propose a reasoning-efficient RL framework that shapes rewards to favor concise yet correct reasoning trajectories. SABER (Zhao et al., 2025) introduces switchable and balanced reward schedules to train models under multiple budget tiers. LSPO (Chen et al., 2025a) develops a length-aware data sampling strategy that dynamically filters training batches based on response length (e.g., retaining the shortest and longest) to stabilize training. Other variants include LASER (Liu et al., 2025b), which introduces adaptive, difficulty-aware reward shaping, penalizing lengthy reasoning more heavily on simple tasks, and Slow-Fast Policy Optimization (SFPO) (Wang et al., 2025g) structures each update into fast, reposition, and slow phases to reduce gradient noise and stabilize convergence. Multi-objective approaches such as BRPO (Qi et al., 2025) optimizes anytime reasoning via budget-sampled dense rewards and a variance-reduced policy gradient, and HBPO (Lyu et al., 2025) structures exploration across hierarchical token-budget groups to preserve capability while reducing average tokens. Together, these works demonstrate how reinforcement learning can encode budget awareness and trade-offs directly into learned reasoning policies through principled reward shaping and constraint design. Other variants include BudgetThinker (Wen et al., 2025) introduces dynamic control tokens whose embeddings encode remaining compute, enabling the model to adjust reasoning depth continuously during decoding, while uniquely combining an RL-based reward function with explicit control tokens, bridging the gap between purely implicit policy optimization and supervised control. AdaReasoner (Wang et al., 2025d) learns an adaptive policy via reinforcement learning to dynamically configure reasoning parameters (e.g., temperature, step length, prompt type) for different tasks
3.1.2 Supervised Fine-tuning and Distillation Methods
A complementary line of work leverages supervised fine-tuning (SFT) or distillation to teach models to produce shorter yet equally valid reasoning chains. C3oT (Kang et al., 2025a) jointly trains on long–short pairs of chain-of-thought (CoT) exemplars, allowing conditional generation of compact rationales without accuracy loss. CoT-Valve (Ma et al., 2025) identifies a controllable “length direction” in parameter space and tunes it with paired long/short data for elastic reasoning compression. TokenSkip (Xia et al., 2025a) prunes semantically redundant tokens and retrains the model under different compression ratios to learn explicit compression control. LS-Mixture SFT (Yu et al., 2025a) integrates long and short CoT data during fine-tuning, enabling adaptive reasoning depth across instances. Distilling 2-1 (Yu et al., 2024) removes explicit reasoning chains during distillation while retaining correctness, encouraging implicit short reasoning behaviors. SoftCoT (Xu et al., 2025b) and SoftCoT++ (Xu et al., 2025c) represent thought processes as continuous latent vectors, allowing smooth trajectory distillation that generalizes across prompts. Skip-Thinking (Chen et al., 2025b) conducts chunk-wise distillation to mitigate gradient saturation in long CoT training. Mixture of Reasonings (MoR) (Xiong et al., 2025) jointly trains LLMs on multiple reasoning strategies (e.g., logical, commonsense, numerical) through multi-style CoT supervision, enabling implicit adaptive reasoning style selection without explicit routing. TALE (Han et al., 2025) performs token-budget–aware prompt tuning that learns to predict minimal feasible token counts per instance. Length-Instruction Fine-Tuning with Chain-of-Thought (LIFT-COT) (Chen and Li, 2025) extends standard supervision with explicit length labels, enabling the model to adhere to user-specified reasoning length or latency constraints while maintaining coherence in its generated rationales. SCoT (Wang et al., 2025b) introduces a speculative chain-of-thought framework where a small draft model rapidly generates multiple reasoning trajectories, and a large target model selects or corrects them, achieving up to 3× faster inference with near-target-level accuracy. Together, these supervised and distillation-based methods internalize compression control within model parameters, achieving adaptive reasoning efficiency without reinforcement objectives.
3.1.3 Learned Controllers and Routers
Another line of work introduces auxiliary controllers or routing modules trained jointly or sequentially with the base model to decide which reasoning path or sub-model to activate. RouteLLM (Ong et al., 2025) learns a lightweight router from preference data to dispatch easy inputs to cheaper small models and difficult ones to larger models, reducing total compute without degrading accuracy. SelfBudgeter (Li et al., 2025b) trains a two-stage system that first predicts the required reasoning budget and then enforces it through gradient-penalized reinforcement optimization. ThinkSwitcher (Liang et al., 2025) extends this idea by training a supervised switching module that decides between concise and elaborate reasoning trajectories based on predicted task difficulty. In Mixture-of-Experts (MoE) models, AdaMOE (Zeng et al., 2024) achieves token-adaptive computation by introducing "null experts" that consume zero FLOPs. A load-balancing loss is used during training to teach the model’s router to adaptively send tokens to either true experts for computation or null experts to skip computation, thereby dynamically allocating resources at the token level. Route-To-Reason (RTR) (Pan et al., 2025) learns to predict accuracy–cost trade-offs across model–strategy pairs and dynamically routes each query to the optimal expert combination under a compute budget. Collectively, these works embed adaptivity at the system-decision level and bridge symbolic routing with learned compute control.
3.2 Training-free Adaptive Reasoning
This section overviews training-free adaptive reasoning, where model parameters remain frozen and adaptivity arises from inference-time control. It introduces three categories: prompt-conditioned, feedback-driven, and modular methods.
Training-free adaptive reasoning denotes inference-time mechanisms that modulate the depth, breadth, or structure of reasoning without any parameter update. Formally, a frozen model $f_{\theta}$ generates a reasoning trajectory $r=(t_{1},\dots,t_{n})$ governed by a control policy $\pi(x,t_{i})$ that dynamically decides continuation or termination based on input difficulty or internal uncertainty. Unlike static efficiency control, these approaches allocate computation adaptively for each instance, emphasizing resource-aware and self-regulated inference (Sui et al., 2025a; Alomrani et al., 2025; Zhu and Li, 2025).
3.2.1 Prompt-conditioned Methods
Explicit token- and instruction-level control. Prompt-conditioned methods introduce explicit instructions, control tokens, or concise templates that determine reasoning length at test time. They embody external adaptivity—computation is modulated through textual conditions rather than learned signals (Cai et al., 2025). Lee et al. (2025) quantifies token complexity to estimate minimal reasoning spans per instance. Contraied-CoT (Nayab et al., 2025) employs templates that prioritize conciseness, and enforce concise reasoning by setting explicit length constraints (e.g., ’100 words’). Sketch-of-Thought (Aytes et al., 2025) introduces cognitive-inspired sketching to generate compact, structured reasoning trajectories. Chain-of-Draft (CoD) (Xu et al., 2025a) introduces a “draft-first” reasoning style: the model writes short, essential reasoning steps (at most  5 words) per turn, thus minimizing verbosity while preserving correctness. ConciseHint (Tang et al., 2025) integrates continuous hints or contextual anchors for difficulty-aware reasoning. Instruction-level prompt schemes such as Plan-and-Budget (Lin et al., 2025) introduce explicit planning and budgeting instructions during inference, enabling the model to allocate reasoning depth and token usage adaptively across sub-questions.
Together, these approaches treat reasoning efficiency as a prompt-engineering problem with explicit, symbolic, yet parameter-free adaptivity.
3.2.2 Feedback-driven Methods
Uncertainty- and entropy-based halting. Feedback-driven adaptation relies on internal feedback such as entropy, confidence, or semantic closure to determine when reasoning should stop or expand. Laaouach (2025) defines entropy-threshold halting, where decoding terminates once uncertainty drops below a pre-defined threshold. Yang et al. (2025) introduces a feedback-driven early exit mechanism that monitors token-level confidence during reasoning to adaptively halt generation once sufficient certainty is reached. Mao et al. (2025) employs a run-jump test to detect answer convergence, terminating the reasoning process when a sufficient number of consecutive steps yield identical intermediate answers. EAGER (Scalena et al., 2025) leverages token-level entropy to branch into multiple reasoning paths only when high uncertainty is detected. Self-Certainty (Kang et al., 2025b) proposes a metric that evaluates answer quality from the model’s own output probabilities, offering a reward-model-free substitute for Best-of-N selection. Sequence-level entropy (Zhu et al., 2025; Sharma and Chopra, 2025) and confidence-based filtering methods such as DeepConf(Fu et al., 2025) further refine this paradigm.
Verifier- and constraint-guided feedback. Beyond internal entropy and confidence, some approaches externalize feedback via verifiers or structural constraints. Liu and Wang (2025) detect answer convergence across partial chains and stop early when consistency stabilizes. Wang et al. (2025a) prune filler thinking tokens by suppressing keywords (e.g., “Wait”, “Hmm”) during decoding. Speculative or constraint-based schemes such as Speculative Search (Wang et al., 2025f) and Value-Guided Search (Wang et al., 2025c) use acceptance thresholds or value models to admit only high-quality thoughts under limited budgets.
MetaReasoner (Sui et al., 2025b) employs a contextual bandit controller that dynamically adjusts reasoning depth during inference based on external validation signals, exemplifying feedback-driven adaptation without reinforcement fine-tuning. These methods remain feedback-driven in essence: reasoning continues or halts based on real-time validation rather than fixed templates.
Consistency- and sampling-based adaptation. A parallel line of work focuses on reasoning breadth—how many rationales to sample before aggregation. Adaptive-Consistency (Aggarwal et al., 2023) and RASC (Wan et al., 2025) estimate consensus and halt when majority agreement stabilizes.
Best-of-NN optimization is improved by Fast BoN (Sun et al., 2024), which employs efficient sampling algorithms to generate high-scoring responses,and ST-BoN (Wang et al., 2025e), which leverages early sampling consistency to identify and truncate suboptimal reasoning paths before full generation, while DPTS (Ding et al., 2025) adaptively prunes search branches. Correctness-first sampling (Li et al., 2025a), also known as Sample Smart, further balance accuracy and latency, refining adaptive breadth control for reasoning LLMs.
Collectively, these methods realize an internal self-regulationthat reasoning is guided by feedback loops rather than pre-specified templates.
3.2.3 Modular Methods
This category covers systems that combine or reuse pre-trained reasoning models through composition, merging, or ensembling. To clarify the boundary with 3.1.3, we treat any static, no-gradient weight interpolation or parameter merging (e.g., long–short or activation-guided merging) as training-free modular adaptation. In contrast, frameworks that involve training new gating, routing, or fusion modules are considered learned controllers or routers (3.1.3).
Pipeline-based adaptive inference. At the system level, modular methods dynamically allocate reasoning compute through a plan–budget–execute pipeline.
Although implemented through decoding-level control rather than explicit architectural modules, AlphaOne (Zhang et al., 2025) employs pacing tokens to regulate cognitive tempo across short-, long-, and tool-augmented reasoning phases. This design mimics modular behavior by dynamically allocating computation among distinct reasoning modes, bridging concise and deliberative thinking within a unified system. CoThink (Fan et al., 2025) introduces a training-free modular pipeline where an instruct model first drafts a concise reasoning outline, and a thinking model expands it, reducing token usage while maintaining accuracy. Industrial frameworks such as Bag-of-Tricks for Inference-Time Compute (Liu et al., 2025a) systematically benchmarks training-free modular techniques that combine generation, verification, and confidence-weighted selection to improve inference-time reasoning efficiency. These systems demonstrate how architectural modularity induces reasoning flexibility without retraining.
Model-merging and ensemble-based adaptation.
A complementary branch combines pre-trained reasoning modules by parameter interpolation or selective ensembling. Long-to-short reasoning via model merging(L2S merging) (Wu et al., 2025) reduces average reasoning length while preserving accuracy, and activation-informed merging(ACM)(Yao et al., 2025) learns layer-wise coefficients from activation mutual information. For language-specific settings, model merging can transfer reasoning ability from a strong “reasoner” into a target-language model without retraining (Pipatanakul et al., 2025). Commercial deployments such as Kimi k1.5 also report hybrid short/long reasoning variants obtained by merging long-CoT and short-CoT specialists (Kimi, 2025). These practices share the same anchor that adaptive compute distribution across fixed modules via weight interpolation or expert ensembling.
4 Conclusion
This survey reframes reasoning in large language models through the lens of adaptivity rather than efficiency alone. It defines adaptivity as the capability to dynamically modulate reasoning depth, structure, or style according to task complexity and uncertainty, provides cognitive formulations of inductive, deductive, and abductive reasoning within LLMs, and introduces a unified taxonomy that distinguishes training-based and training-free implementations. Together, these contributions establish a coherent framework for analyzing and advancing adaptive reasoning toward more efficient, flexible, and context-sensitive language models.
5 Limitations
This survey inevitably faces two limitations. First, due to space constraints, we focus on representative adaptive reasoning methods and do not exhaustively cover all recent variants, especially those emerging in multimodal or agentic settings. Our taxonomy emphasizes conceptual clarity over completeness, which may overlook some niche but relevant techniques. Second, the field of adaptive reasoning is evolving rapidly and new paradigms such as self-improving reasoning, adaptive reflection, and meta-evaluation appear almost monthly. Consequently, future developments may shift or refine the taxonomy presented here.